OpenStudy (anonymous):

Solve the differential equation dx+(x/y-sin(y))dy=0.

OpenStudy (anonymous):

Dividing by dx and simplifying we get,\[y \sin y \times dy = x dy + y dx\] \[y \sin y \times dy = d(xy)\] exact differential Now integrate on both sides by using integration by parts.

OpenStudy (loser66):

\[(\frac {x}{y}-sin y)dy =-dx\] integral both sides xlny +cos y = -x +C xlny +cos y +x =C is it not right? I am new in differential equation. Please, check

OpenStudy (anonymous):

No loser66, you cannot integrate both the dependent and independent variable simultaneously.

OpenStudy (loser66):

why? the LHS respect to y, and the RHS respect to x.

OpenStudy (anonymous):

I am talking about x/y.

OpenStudy (loser66):

oh yea, thank you. So, we must have x and y separately when calculating, right?

OpenStudy (anonymous):

The correct answer is xy+ycos(y)-sin(y)=c but you need to use integrating factor for this problem.

OpenStudy (anonymous):

Yes, thats it. Differential equations is a good topic. Try learning it.

OpenStudy (anonymous):

Is it compulsory to use integrating factor for this problem? I got the answer in my way.

OpenStudy (anonymous):

Thanks, guys.